OpenStudy (anonymous):

factor: 3xsquared - 54x = 243

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

well, you know its 3x and x

OpenStudy (anonymous):

so what multiplies by 3 and adds x to get -54?

OpenStudy (anonymous):

x=9(1-sqrt2) x=9)1+sqrt2)

OpenStudy (anonymous):

giggles they dont want to solve it, they want to factor it.

OpenStudy (anonymous):

(3x)^2(sqrt-54x)^2=243^2 9x^2-54=59049 x^2-54x-6561=0 use quadratic equation to find x

hero (hero):

That's not factoring, but okay

OpenStudy (anonymous):

I don't think it'll factor into anything that looks nice.

OpenStudy (anonymous):

I agree with TBates. Usually C is not as large as 243.

hero (hero):

Any quadratic is factorable. I didn't say anything about it looking nice though

OpenStudy (anonymous):

K we understand that Hero we are just commenting about the extensiveness of the quadratic. Thats all.

hero (hero):

I factored it and got: \[(3x - 27\sqrt{2} - 27) (x + 9\sqrt{2} - 9)\]

OpenStudy (anonymous):

x=-31.38 or x=139.38

hero (hero):

If you want to know how I got that, just ask

OpenStudy (anonymous):

how did you get that?

hero (hero):

lol

hero (hero):

I factored by grouping

hero (hero):

Okay, I'll show you what I did...

OpenStudy (anonymous):

why thank you hero.

hero (hero):

Okay, so this is the formula I used:  You input the coefficients of a, b, and c into the formula and solve for d. (You probably didn't know a "d" existed...but it does..)